OpenStudy (anonymous):

Find the range of the function y = -5x - 7

OpenStudy (hockeychick23):

You need to plug in x values first so lets just choose a few numbers. x=0,-1,-2,-3 y= -5(0)-7 y=-7 y=-5(-1)-7 y=-2 y=-5(-2)-7 y=3 y=-5(-3)-7 y=8 domain or x values: 0,-1,-2, -3 range or y values: -7,-2, 3, 8

OpenStudy (anonymous):

The range is all real numbers, you can get any value for y from (-infinity, infinity)

OpenStudy (anonymous):

as the output

OpenStudy (anonymous):

Thank you!!